# 2]organization=Huadian Coal Industry Group Co. Ltd.,country=China

### v1

@@ Introduction
$+u_{12}x_{1}x_{2}+u_{13}x_{1}x_{3}+u_{23}x_{2}x_{3}$ $\mathbf{H_{5}}=u_{11}+u_{22}+u_{12}$ $\mathbf{H_{6}}=u_{11}+u_{33}+u_{13}$ $\mathbf{H_{7}}=u_{22}+u_{33}+u_{23}$ $\mathbf{H_{8}}=0$
@@ 3.2 Neural Network Compression
Figure 3: (a) Qubits in yellow are mapped from the pruning parameters set $\mathcal{P}$, and (b) qubits in blue are mapped from the quantization parameters set $\mathcal{Q}$. In this example, we assume a model with three layers, containing $3$, $2$, and $4$ filters respectively, with weights represented by 8 bits.
@@ 4.1 Mapping of Pruning Parameters
Here, $\textbf{w}_{n,i}$ represents the group of weights in filter $i$ of layer $n$, while $p_{n,i}$ denotes the corresponding mask applied to that filter. Each parameter $p_{n,i}$ can take a binary value, either 00 or $1$. Specifically, $p_{n,i}=1$ indicates that the associated filter is pruned, while $p_{n,i}=0$ signifies that the filter is retained.
@@ 4.2 Mapping of Quantization Parameters
$p_{n,i}$ [Uncaptioned image] $D_{n,i}$
$q_{n,k}$ [Uncaptioned image] $E_{n,k}$
$p_{n,i}$ $q_{n,k}$ [Uncaptioned image] $-F_{n,i,k}$
@@
$p_{n,i}$ [Uncaptioned image] $A_{n,i,i}-\gamma\cdot{D}_{n,i}$
$q_{n,k}$ [Uncaptioned image] $\beta\cdot B_{n,k,k}-\gamma\cdot{E}_{n,k}$
@@ 5.2 Quantization Loss
To convert the formulation of RMSE into QUBO form, we take the logarithm of RMSE$/c_{n}$ and square the result to obtain the quantization loss component, i.e., $L_{q}(\mathcal{Q})$, which is expressed in Equation (5.2). The coefficients of two qubits, $q_{n,k}$ and $q_{n,l}$, are represented by $B_{n,k,l}$ and can be pre-calculated before decoding quantum optimization. Accordingly, the coupling map of qubits is shown in Table 3.
@@ 5.4 Combined QUBO Model
Refer to Algorithm 1, we begin with an initial $\beta$ value of ${\|{\mathbf{A}}\|_{1}}/{\|\mathbf{B}\|_{1}}$ and an arbitrary $\gamma$ value (line 1). As coefficients ${A}$ and ${B}$ approximate the severity of pruning loss and quantization error respectively, we hypothesize that the $l1$-norm of these coefficients across the model should be similar in magnitude to balance between pruning and quantization. From our experiments, we find that this initial $\beta$ hypothesis is generally near-optimal, providing a useful starting point for the search.
@@
    Fixing $\beta$, determine an appropriate range of $\gamma$ values in which to search (lines 6-12). Specifically, we aim to search in the range of $[\gamma_{l},\,\gamma_{u}]$, for which $\gamma_{l}$ has high accuracy above threshold $a_{th}$ while $\gamma_{u}$ has accuracy below threshold $a_{th}$
@@
    Perform binary search for the maximum $\gamma$ in the range $[\gamma_{l},\,\gamma_{u}]$ that fulfills the accuracy threshold (line 13). Let this value be $\gamma_{n}$
@@ 7.3 Compression Performance by AQC
Figure 6: Comparison of quantum annealing (QA), genetic algorithms (GA), reinforcement learning (RL), and machine learning package tools (PT) regarding accuracy, model reduction ratio, and optimization time. We implemented the GA algorithms (NSGA-II) using code from Reference [25] and QA algorithms (SAC) using code from Reference [24], tuning hyperparameters until the optimization converged. For the PT method, we utilize quantization aware training (QAT) [18] and Ln-structured pruning [17] algorithms, both integrated into the PyTorch framework.

### v2

@@ Introduction
$u_{12}x_{1}x_{2}+u_{13}x_{1}x_{3}+u_{23}x_{2}x_{3}$ $\mathbf{H_{5}}=u_{11}+u_{22}+u_{12}$ $\mathbf{H_{6}}=u_{11}+u_{33}+u_{13}$ $\mathbf{H_{7}}=u_{22}+u_{33}+u_{23}$ $\mathbf{H_{8}}=0$
@@ 3.2 Neural Network Compression
Figure 3: (a) Qubits in yellow are mapped from the pruning parameters set $\mathcal{P}$, and (b) qubits in blue are mapped from the quantization parameters set $\mathcal{Q}$. In this example, we assume a model with three layers, containing 33, 22, and 44 filters respectively, with weights represented by 8 bits.
@@ 4.1 Mapping of Pruning Parameters
Here, $\textbf{w}_{n,i}$ represents the group of weights in filter $i$ of layer $n$, while $p_{n,i}$ denotes the corresponding mask applied to that filter. Each parameter $p_{n,i}$ can take a binary value, either 0 or $1$. Specifically, $p_{n,i}=1$ indicates that the associated filter is pruned, while $p_{n,i}=0$ signifies that the filter is retained.
@@ 4.2 Mapping of Quantization Parameters
$p_{n,i}$ - [Uncaptioned image] $D_{n,i}$
$q_{n,k}$ - [Uncaptioned image] $E_{n,k}$
$p_{n,i}$ $q_{n,k}$ [Uncaptioned image] $F_{n,i,k}$
@@
$p_{n,i}$ - [Uncaptioned image] $A_{n,i,i}-\gamma\cdot{D}_{n,i}$
$q_{n,k}$ - [Uncaptioned image] $\beta\cdot B_{n,k,k}-\gamma\cdot{E}_{n,k}$
@@ 5.2 Quantization Loss
[…]
@@ 5.4 Combined QUBO Model
Refer to Algorithm 1, we begin with an initial $\beta$ value of ${\|{\mathbf{A}}\|_{1}}/{\|\mathbf{B}\|_{1}}$ and an arbitrary $\gamma$ value (line 1). As coefficients A{A} and B{B} approximate the severity of pruning loss and quantization error respectively, we hypothesize that the $l1$-norm of these coefficients across the model should be similar in magnitude to balance between pruning and quantization. From our experiments, we find that this initial $\beta$ hypothesis is generally near-optimal, providing a useful starting point for the search.
@@
    Fixing $\beta$, determine an appropriate range of $\gamma$ values in which to search (lines 6-12). Specifically, we aim to search in the range of $\gamma_{l},\,\gamma_{u}]$, for which $\gamma_{l}$ has high accuracy above threshold $a_{th}$ while $\gamma_{u}$ has accuracy below threshold $a_{th}$ .
@@
    Perform binary search for the maximum $\gamma$ in the range $\gamma_{l},\,\gamma_{u}]$ that fulfills the accuracy threshold (line 13). Let this value be $\gamma_{n}$
@@ 7.3 Compression Performance by AQC
Figure 6: Comparison of quantum annealing (QA), genetic algorithms (GA), reinforcement learning (RL), and machine learning package tools (PT) regarding accuracy, model reduction ratio, and optimization time. We implemented the GA algorithms (NSGA-II) using code from Reference [25] and RL algorithms (SAC) using code from Reference [24], tuning hyperparameters until the optimization converged. For the PT method, we utilize quantization aware training (QAT) [18] and Ln-structured pruning [17] algorithms, both integrated into the PyTorch framework.